# Weyl symmetry without the traceless condition

### v3

@@ VII Weyl invariant cosmology and the many-worlds perspective
The CGR theory (37) is a much more versatile alternative to solving current gravitational puzzles than GR alone, since the latter belongs to the class of conformal equivalence of CGR. Suppose that we have an GR solution $\hat{g}_{\mu\nu}(x)$ and that a gravitational state is labeled “kk”;
${\cal S}_{k}:\left(\phi_{k}(x),g^{(k)}_{\mu\nu}(x),...\right),\;\;k=1,2,...\,,N,$ (64)
@@
where $N\rightarrow\infty$ and the ellipsis stands for other fields coupled to gravity, is such that the fields $\phi_{k}(x)$ and $g^{(k)}_{\mu\nu}(x)$ satisfy the CGR-EOM (38) for any $k$ and any gauge choice $\phi_{k}=\phi_{k}(x)$. The different metric tensors $g^{(k)}_{\mu\nu}(x)$ are related to the GR solution through (52);
$g^{(k)}_{\mu\nu}(x)=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}(x)}\right)^{2}\hat{g}_{\mu\nu}(x).$
@@
Hence, instead of a single GR gravitational state ${\cal S}_{\text{gr}}$, in CGR there is an infinite set of gravitational states: $\mathfrak{S}_{\text{cgr}}=\{{\cal S}_{\text{gr}},{\cal S}_{1},{\cal S}_{2},...,{\cal S}_{N}\}$, including the GR gauge itself, which are available to describe a given set of gravitational phenomena.
@@
where $t$ is cosmic time, $a(t)$ is the scale factor, and the latin indices run through space: $i,j=1,2,3$. We further assume that the background matter content is a perfect fluid with energy density $\rho$ and pressure $p$ that satisfy the following EOS: $p=(\gamma-1)\rho$, where $\gamma$ is the barotropic index of the fluid. The independent equations are the $(0,0)$-component of the Einstein-type EOM (38) and the nonhomogeneous continuity equation (39);
@@
In CGR theory, instead of a single gravitational state for each value of the barotropic index $\gamma$, as in GR, we have an infinite class of gravitational states:
$\mathfrak{S}_{\text{cgr}}=\left\{\left(\sqrt{6}M_{\text{pl}},a_{\text{gr}}(\tau)\right),\left(\phi_{1}(\tau),a_{1}(\tau)\right),...,(\phi_{N}(\tau),a_{N}(\tau))\right\}.$
@@
$a_{k}(\tau)=\left[\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}(\tau)}\right]a_{\text{gr}}(\tau),\;\;k=1,2,...,N.$
The large (in principle infinite) variety of cosmic behaviors in $\mathfrak{S}_{\text{cgr}}$ is what makes the CGR theory and, therefore, Weyl symmetry attractive within the cosmological context and, at the same time, yields the most unappealing feature of this theory when it is analyzed under the active approach: the loss of predictability. The Weyl invariant CGR theory is not capable of defining a unique gravitational state, but a whole class of them, so that the definition of the state that correctly describes a given gravitational phenomenon is left to experiment. Nevertheless, this theory can be assigned some predictive power. For example, the fifth force $f^{\mu}=h^{\mu\lambda}\partial_{\lambda}\phi/\phi$ in (40), which is associated with point-dependent masses so it can pass unnoticed in local experiments, points to a modification of the redshift measurements that could explain, in principle, the present accelerated pace of the cosmic expansion and so-called dark matter, because it does not interact with radiation ($T^{\text{rad}}=0$ $\Rightarrow$ $f_{\mu}=0$). In the last instance, the active approach to Weyl symmetry within CGR theory may be viewed as a classical version of the many-worlds interpretation of gravitational interactions of matter everett_rmp_1957 ; wheeler ; dewitt_phys_rev_1967 ; dewitt ; barvinsky_cqg_1990 ; omnes_rmp_1992 ; tegmark_1998 ; garriga_prd_2001 ; zurek_rmp_2003 ; tegmark_nature_2007 ; quiros_prd_2023 .
@@
Additionally, many measurements in cosmology rest on the emission and detection of photons. These follow null-geodesics of Riemann geometry, which are themselves form-invariant under WTs. Any measuring clock or stick, which is made of a combination of emitted and/or absorbed photons, is not transformed by the Weyl transformations and so is the related measuring unit. Hence, the above argument faraoni_prd_2007 is not universal and can be circumvented. In addition, even if a measuring unit, for example, the mass unit, is point-dependent, there are ways in which this effect can be experimentally revealed. For example, the point-dependent property of the masses of timelike fields, and so the fifth force appearing in (39), cannot be detected in local experiments because any measuring stick is made up of fields, so it suffers the same universal point-dependence as the mass being measured, and the effect cancels out. However, in experiments that involve the detection of photons emitted at a distant point in spacetime, namely redshift measurements, the effect can be measured (at least in principle).
@@ VIII Observational consequences of CGR: The redshift measurements
Let us explain the peculiar contribution to the redshift of frequencies carried by the Weyl symmetry. In CGR theory, two phenomena contribute to the redshift of frequencies: 1) the gravitational redshift (also curvature redshift) and 2) the redshift due to variation of masses in spacetime (mass redshift). The former effect is due to the propagation of photons in a curved background, while the latter is due to the point-dependent property of masses: $m=m(x)=\mu\,\phi(x)$. This mass dependence affects the redshift because the frequency of a photon emitted/absorbed by some atom (atomic number $Z$) is given by the Rydberg formula:
@@
where $m_{e}$ is the mass of the electron, $c$ is the speed of light, $h$ is the Planck constant (here, temporarily we assume that $c\neq h\neq 1$), $\alpha$ is the fine-structure constant ($\alpha\approx 1/137$), while $n_{l}$ and $n_{u}$ are the principal quantum numbers of the lower and upper energy levels, respectively. Suppose that two identical atoms are located at spacetime points $x$ and $x_{0}$, respectively. The relative shift of the frequency of the emitted/absorbed photons is given by
$z_{m}=\frac{\gamma_{e/a}(x)-\gamma_{e/a}(x_{0})}{\gamma_{e/a}(x_{0})}=\frac{m_{e}(x)-m_{e}(x_{0})}{m_{e}(x_{0})}=\frac{\phi(x)}{\phi(x_{0})}-1.$ (72)
Note that this emission/absorption frequency does not depend on the specific atom that emits/absorbs the photon, nor on the specific atomic transition $n_{u}\rightleftharpoons n_{l}$. Note also that in the GR gauge, since $\phi(x)=\phi(x_{0})=M_{\text{pl}}$, this type of redshift does not arise. Meanwhile, the curvature redshift is due to the dynamics of the photon frequency during its propagation in a curved background. The frequency of a photon emitted/absorbed at some point $x$ in spacetime can be found as a solution of the null-geodesics (42): $\gamma_{e/a}(x)=\gamma^{0}_{e/a}/\sqrt{-g_{00}(x)}$, where $\gamma^{0}_{e/a}$ is some constant frequency. The curvature redshift resulting from the propagation of a photon from point $x$ to $x_{0}$ and the overall redshift are given by:
$z_{c}=\frac{\gamma_{e/a}(x)-\gamma_{e/a}(x_{0})}{\gamma_{e/a}(x_{0})}=\sqrt{\frac{g_{00}(x_{0})}{g_{00}(x)}}-1,$ (73)
@@
$z_{\text{tot}}=z_{c}+z_{m}=\sqrt{\frac{g_{00}(x_{0})}{g_{00}(x)}}+\frac{\phi(x)}{\phi(x_{0})}-2,$ (74)
respectively. If we take into account the relationship (52) between the metric $g_{\mu\nu}(x)$ of an arbitrary gauge and the metric $\hat{g}_{\mu\nu}(x)$ in the GR gauge, we get,
$z_{\text{tot}}=\frac{\phi(x)}{\phi(x_{0})}\sqrt{\frac{\hat{g}_{00}(x_{0})}{\hat{g}_{00}(x)}}+\frac{\phi(x)}{\phi(x_{0})}-2,$
or if we realize that the GR redshift $z=\sqrt{\hat{g}_{00}(x_{0})/\hat{g}_{00}(x)}-1$, then
$z_{\text{tot}}=\frac{\phi(x)}{\phi(x_{0})}\,z+2\left(\frac{\phi(x)}{\phi(x_{0})}-1\right).$ (75)
In a cosmological context, if we replace $t\rightarrow x$, where $t$ is cosmic time, and set $x_{0}=0$ (present time), under the assumption of very small $\phi$-variations, we can write
$\frac{\phi(t)}{\phi(0)}=1+\frac{\dot{\phi}(0)}{\phi(0)}\Delta t=1+\epsilon,$ (76)
@@
where a dot means derivative with respect to cosmic time, we assume that $\phi(0)=\sqrt{6}M_{\text{pl}}$, and that $\epsilon>0$ is a small positive number, so that
$z_{\text{tot}}=(1+\epsilon)z+2\epsilon.$ (77)
@@
This simplified expression shows that, in CGR cosmology, the measured redshift $z_{\text{tot}}$ is greater than the redshift according to GR. Although (77) is an approximate equation, for redshifts $\sim 1$ and a bit larger, this result can have obvious implications for the explanation of the late-time dynamics of our Universe. Since the fifth force that arises in (39) and in (40) is due to the point-dependent property of the mass: $m(x)=\kappa\,\phi(x)$, accurate measurements of the second term in the RHS of (77) would entail indirect measurements of the extra force $f^{\alpha}=h^{\alpha\lambda}\partial_{\lambda}\phi/\phi$.

### v4

@@ VII Weyl-invariant cosmology and the many-worlds perspective
The CGR theory (24) is a much more versatile alternative to solving current gravitational puzzles than GR alone, since the latter belongs to the class of conformal equivalence class $\mathfrak{C}_{\text{conf}}$ of CGR. Suppose that we have an arbitrary GR solution labeled “AA” (from now on the upper case latin indexes $A$, $B$, $C$, etc., run over all possible solutions of GR-EOM $A=1,2,3,...,N_{s}$, where $N_{s}$ is the number of different solutions), and that a corresponding global gravitational state labeled “kk” is represented as (compare with (50));
$\mathfrak{G}(A|k):\{\phi_{k}(x),g^{(A|k)}_{\mu\nu}(x),...\},\;\;A=1,2,...,N_{s};\;\;k=1,2,...,N,$ (63)
@@
where $N\rightarrow\infty$, and the ellipsis in curly brackets stands for other fields coupled to gravity. Let us assume that the fields $\phi_{k}(x)$ and $g^{(A|k)}_{\mu\nu}(x)$ satisfy the CGR-EOM (35) for any $k$ and any gauge choice $\phi_{k}=\phi_{k}(x)$. The different metric tensors $g^{(A|k)}_{\mu\nu}(x)$ are related to the GR solution labeled “AA” through (49);
$g^{(A|k)}_{\mu\nu}(x)=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}(x)}\right)^{2}\hat{g}^{(A)}_{\mu\nu}(x),$
@@
where $\hat{g}^{(1)}_{\mu\nu}(x)$, $\hat{g}^{(2)}_{\mu\nu}(x)$$\hat{g}^{(N_{s})}_{\mu\nu}(x)$, are different solutions of the GR-EOM. Hence, instead of a single GGS: $\mathfrak{G}_{\text{gr}}(A):\{\sqrt{6}M_{\text{pl}},\hat{g}^{(A)}_{\mu\nu}(x),$...\}, corresponding to each GR solution “AA”, in CGR there is an infinite set of global gravitational states: $\mathfrak{C}_{\text{conf}}(A)=\{\mathfrak{G}_{\text{gr}}(A),\mathfrak{G}(A|1),\mathfrak{G}(A|2),...,\mathfrak{G}(A|N-1)\}$ ($A=1,2,...,N_{s}$), including the GR solution itself, which are available to describe a given set of gravitational phenomena. Only experiments can pick out the GGS in $\mathfrak{C}_{\text{conf}}(A)$ that more closely describes given phenomena.
@@
where $t$ is cosmic time, $a(t)$ is the scale factor, and the lower case latin indices run through space: $i,j=1,2,3$. We further assume that the background matter content is a perfect fluid with energy density $\rho$ and pressure $p$ that satisfy the following EOS: $p=(\gamma-1)\rho$, where $\gamma$ is the barotropic index of the fluid. The independent equations are the $(0,0)$-component of the Einstein-type EOM (35) and the nonhomogeneous continuity equation (36);
@@
In CGR theory, instead of a single GGS for each value of the barotropic index $\gamma$, as in GR, we have an infinite class of global gravitational states:
$\mathfrak{C}_{\text{conf}}=\left\{\left(\sqrt{6}M_{\text{pl}},a_{\text{gr}}(\tau)\right),\left(\phi_{1}(\tau),a_{1}(\tau)\right),...,(\phi_{N-1}(\tau),a_{N-1}(\tau))\right\}.$
@@
$a_{k}(\tau)=\left[\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}(\tau)}\right]a_{\text{gr}}(\tau),\;\;k=0,1,2,...,N-1,$
where the label $k=0$ is used exclusively to denote the GR gauge. That is, $\phi_{0}=\sqrt{6}M_{\text{pl}}$.
The large (in principle infinite) variety of cosmic behaviors in $\mathfrak{S}_{\text{cgr}}$ is what makes the CGR theory and, therefore, Weyl symmetry attractive within the cosmological context and, at the same time, yields the most unappealing feature of this theory when it is analyzed under the active approach: the loss of predictability. The Weyl-invariant CGR theory is not capable of defining a unique GGS, but a whole class of them, so that the definition of the state that correctly describes a given gravitational phenomenon is left to experiment. Nevertheless, this theory can be assigned some predictive power. For example, the fifth force $f^{\mu}=h^{\mu\lambda}\partial_{\lambda}\phi/\phi$ in (37), which is associated with point-dependent masses so it can pass unnoticed in local experiments, points to a modification of the redshift measurements that could explain, in principle, the present accelerated pace of the cosmic expansion and so-called dark matter, because it does not interact with radiation ($T^{\text{rad}}=0$ $\Rightarrow$ $f_{\mu}=0$). In the last instance, the active approach to Weyl symmetry within CGR theory may be viewed as a classical version of the many-worlds interpretation of gravitational interactions of matter everett_rmp_1957 ; wheeler ; dewitt_phys_rev_1967 ; dewitt ; barvinsky_cqg_1990 ; omnes_rmp_1992 ; tegmark_1998 ; garriga_prd_2001 ; zurek_rmp_2003 ; tegmark_nature_2007 ; quiros_prd_2023 .
@@
Additionally, many measurements in cosmology rest on the emission and detection of photons. These follow null-geodesics of Riemann geometry, which are themselves form-invariant under WTs. Any measuring clock or stick, which is made of a combination of emitted and/or absorbed photons, is not transformed by the Weyl transformations and so is the related measuring unit. Hence, the above argument faraoni_prd_2007 is not universal and can be circumvented.
@@ VIII Observational consequences of CGR: The redshift measurements
Even if measuring units, for example, the mass unit, are point-dependent, there are ways in which this effect can be experimentally revealed. For example, the point-dependent property of the masses of timelike fields, and so the fifth force appearing in (36), cannot be detected in local experiments because any measuring stick is made up of fields, so it suffers the same universal point-dependence as the mass being measured, and the effect cancels out. However, in experiments that involve the detection of photons emitted at a distant point in spacetime, namely redshift measurements, at least in principle, the effect can be measured.
Let us explain the peculiar contribution to the redshift of frequencies carried by the Weyl symmetry. In CGR theory, two phenomena contribute to the redshift of frequencies: 1) the gravitational redshift (also curvature redshift) and 2) the redshift due to variation of masses in spacetime (mass redshift). The former effect is due to the propagation of photons in a curved background, while the latter is due to the point-dependent property of masses: $m=m(x)=\kappa\,\phi(x)$. This mass dependence affects the redshift because the frequency of a photon emitted/absorbed by some atom (atomic number $Z$) is given by the Rydberg formula:
@@
where $m_{e}$ is the mass of the electron, $c$ is the speed of light, $h$ is the Planck constant (here, temporarily we assume that $c\neq h\neq 1$), $\alpha$ is the fine-structure constant ($\alpha\approx 1/137$), while $n_{l}$ and $n_{u}$ are the principal quantum numbers of the lower and upper energy levels, respectively. Consider two identical atoms, one of which emits a photon at some time $t$ in the past, and the other atom absorbs the photon at present time $t_{0}$. The relative shift of the frequency of the emitted/absorbed photons is given by
$z_{m}=\frac{\gamma_{e/a}(t)-\gamma_{e/a}(t_{0})}{\gamma_{e/a}(t_{0})}=\frac{m_{e}(t)-m_{e}(t_{0})}{m_{e}(t_{0})}=\frac{\phi(t)}{\phi(t_{0})}-1.$ (71)
Note that this emission/absorption frequency does not depend on the specific atom that emits/absorbs the photon, nor on the specific atomic transition $n_{u}\rightleftharpoons n_{l}$. Note also that in the GR gauge, since $\phi(t)=\phi(t_{0})=\sqrt{6}M_{\text{pl}}$, this type of redshift does not arise. Meanwhile, the curvature redshift is due to the dynamics of the photon frequency during its propagation in a curved background. The frequency of a photon at any time $t$ while it propagates in an expanding universe can be found as a solution of the null-geodesics (39) (see equation (101) of appendix A): $\gamma(t)=\gamma_{0}/a(t)$, where $\gamma_{0}$ is some constant frequency. The curvature redshift resulting from the propagation of a photon in an expanding (homogeneous and isotropic) FRW background between times $t$ and $t_{0}$, and the overall redshift are given by:
$z_{c}=\frac{\gamma(x)-\gamma(x_{0})}{\gamma(x_{0})}=\frac{a(t_{0})}{a(t)}-1\;\Rightarrow\;a(t)=\frac{a(t_{0})}{1+z_{c}},$ (72)
@@
$z_{\text{tot}}=z_{c}+z_{m}=\frac{a(t_{0})}{a(t)}+\frac{\phi(t)}{\phi(t_{0})}-2,$ (73)
respectively. If we take into account the relationship (46); that is,
$a(t)=\frac{\sqrt{6}M_{\text{pl}}}{\phi(t)}\,\hat{a}(t),$
where $\hat{a}$ is the scale factor in the GR gauge, we get that
$z_{\text{tot}}=\frac{\phi(t)}{\phi(t_{0})}\frac{\hat{a}(t_{0})}{\hat{a}(t)}+\frac{\phi(t)}{\phi(t_{0})}-2,$
or if we realize that the GR redshift $z={\hat{a}(t_{0})/\hat{a}(t)}-1$, then
$z_{\text{tot}}=\frac{\phi(t)}{\phi(t_{0})}\,z+2\left(\frac{\phi(t)}{\phi(t_{0})}-1\right).$ (74)
In a cosmological context, we replace $t\rightarrow z$, and set $t_{0}=0$ (present time):
$z_{\text{tot}}=\frac{\phi(z)}{\phi(0)}\,z+2\left(\frac{\phi(z)}{\phi(0)}-1\right).$ (75)
Let us for simplicity and definiteness choose the linear $z$ gauge
$\frac{\phi(z)}{\phi(0)}=1+\frac{\phi^{\prime}(0)}{\phi(0)}z=1+\mu z,$ (76)
@@
where a prime means derivative with respect to $z$ and $\mu$ is a dimensionless constant. We assume that $\phi(0)=\sqrt{6}M_{\text{pl}}$. Hence, in this “linear gauge”
$z_{\text{tot}}=(1+2\mu+\mu z)z.$ (77)
@@
The expression (77) shows that in this gauge of CGR cosmology, the measured redshift $z_{\text{tot}}$ is greater than the redshift according to the GR gauge. This result can have obvious implications for the explanation of the late-time dynamics of our Universe. Since the fifth force that arises in (36) and in (37) is due to the point-dependent property of the mass: $m(x)=\kappa\,\phi(x)$, accurate measurements of the second term in the RHS of (75); $2\mu z$, would entail indirect measurements of the extra force $f^{\alpha}=h^{\alpha\lambda}\partial_{\lambda}\phi/\phi$.
VIII.1 SNIa redshift measurements: Weyl symmetry vs dark energy
[…]
Measuring distances to distant stars is a challenging task. Several quantities, such as the apparent magnitude $m$, the absolute magnitude $M$ (logarithmic measures of flux and luminosity, respectively), and the luminosity distance $d_{L}$ are involved ccp-padman ; dodelson-book . The latter is related to the energy flux $\Phi$ measured by the observer at $z=0$, which comes from a distant source with actual luminosity $L$; $d_{L}=\sqrt{L/4\pi\Phi}$. It is, in general, a model-dependent quantity. It can be related to theoretical quantities as it follows dodelson-book :
$d_{L}=\frac{d_{C}}{a}=(1+z)d_{C},$ (78)
where the comoving distance $d_{C}$ is given by
$d_{C}=\int_{t}^{0}\frac{dt^{\prime}}{a(t^{\prime})}=\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})}.$ (79)
[…]
$H(z)=H_{0}\sqrt{\Omega_{m,0}(z+1)^{3}+\Omega_{\Lambda,0}},$ (80)
where the dimensionless energy density of matter and of the cosmological constant are given by:
$\Omega_{m,0}\equiv\frac{\rho_{0}}{3H^{2}_{0}},\;\Omega_{\Lambda,0}\equiv\frac{\Lambda}{3H^{2}_{0}},$
respectively. In the above equations, $H_{0}$ and $\rho_{m,0}$ are the present values of the Hubble parameter and of the matter density.
For simplicity, we shall investigate the modification to the redshift in the linear gauge (76) where the relationship (77) between the overall (measured) redshift in this gauge and the GR redshift $z$ takes place. Note that in the expression for the comoving distance $d_{C}$, the expression under the integral: $dt/a(t)$ is a Weyl form-invariant quantity, so it has the same functional dependence in the linear gauge and in the GR gauge ($\Lambda=0$, $\Omega_{m,0}=1$). That is, $\int_{t}^{0}dt^{\prime}/a(t^{\prime})=\int_{0}^{z}dz^{\prime}/(z^{\prime}+1)^{3/2}$. However, a subtlety must be considered: while the integration limit in GR gauge is $z$, in the “linear gauge” it must be $z_{\text{tot}}$, which is the measured redshift in that gauge. This means that in GR gauge we have:
$H_{0}d_{C}=\int_{0}^{z}\frac{dz^{\prime}}{(z^{\prime}+1)^{3/2}}=2\left(\frac{\sqrt{z+1}-1}{\sqrt{z+1}}\right),$ (81)
while in the linear gauge (77) we get that
$H_{0}d^{\text{lin}}_{C}=\int_{0}^{z_{\text{tot}}}\frac{dz^{\prime}}{(z^{\prime}+1)^{3/2}}=2\left(\frac{\sqrt{z_{\text{tot}}+1}-1}{\sqrt{z_{\text{tot}}+1}}\right)=2\left(1-\frac{1}{\sqrt{1+(1+2\mu+\mu z)z}}\right),$ (82)
where $z_{\text{tot}}$ is given by (77).
Another subtlety to be considered is that, according to (72) and (78), in an arbitrary gauge $d_{L}=d_{C}/a=d_{C}(1+z_{c})$. That is, only the curvature redshift matters, since it is the one related to the scale factor (72) (see EQ. (102) in appendix A). Hence, in the linear gauge (76), the expression for the luminosity distance is given by:
$H_{0}d^{\text{lin}}_{L}=(1+z_{c})H_{0}d^{\text{lin}}_{C}=(1+\mu z)(1+z)H_{0}d^{\text{lin}}_{C},$
where $H_{0}d^{\text{lin}}_{C}$ is given by (82).
According to Eq. (31) of reference ccp-padman , the redshift-dependent apparent magnitude $m$ of distant SNIa is given by [^20] Since we consider the dimensionless combination $H_{0}d_{L}$ rather than $d_{L}$, this means that in the equation to determine the apparent magnitude a term $5\log_{10}h$ is to be taken away, where $H_{0}=100h\,\text{km}\,\text{s}^{-1}\,\text{Mpc}^{-1}$.
$m=M-5\log_{10}h+42.38+5\log_{10}\left(H_{0}d_{L}\right).$
Here we take M=19.09M=-19.09 and $h=0.72$ so that,
$m=5\log_{10}(H_{0}d_{L})+24,$ (83)
which is the master equation to determine the theoretical values of the apparent magnitude.
[…]
Now we are in a position to explain how the accelerated expansion is naturally avoided in the present conformal invariant theory. As an illustration, we shall consider one of the first data sets on high redshift supernovae measurements perlmutter-1999 , which provided early evidence for accelerated expansion of the universe. We shall consider three cases: 1) the GR model, where
$H_{0}d_{L}=(1+z)H_{0}d_{C},$ (84)
[…]
$H_{0}d^{\Lambda}_{L}=(1+z)H_{0}d^{\Lambda}_{C},\;H_{0}d^{\Lambda}_{C}=\int_{0}^{z}\frac{dz^{\prime}}{\sqrt{\Omega_{m,0}(z^{\prime}+1)^{3}+\Omega_{\Lambda,0}}},$ (85)
and 3) the Weyl-invariant CGR model:
$H_{0}d^{\text{lin}}_{L}=(1+\mu z)(1+z)H_{0}d^{\text{lin}}_{C},$ (86)
where $H_{0}d^{\text{lin}}_{C}$ is given by (82). We compute the theoretical values of the apparent magnitude $m$ (83) for the three cases and compare the results. These are checked against redshift data.
[…]
VIII.2 Pound-Rebka-Snider experiments